# Reconstrucion of oriented matroids from Varchenko-Gelfand algebras

### v1

@@ 1. Introduction
(4.3) Let $R$ be an integral domain with $\operatorname{char}R\neq 2$, and let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be real arrangements. If $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})_{R}\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})_{R}$ as graded $R$-algebras, then the oriented matroids of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are isomorphic.
@@
These results show that VG algebras contain sufficient information to recover oriented matroids under genericity conditions. We also point out that none of the following invariants determine the VG algebra: the intersection lattice $L(\mathcal{A})$, the Orlik-Solomon algebra $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A})$, or the Poincaré polynomial of $M(\mathcal{A})$ (Example 3.10, Example 5.2, Example 5.4). These relationships are summarized in Figure 1. (See also Section 2.4 for details.)
@@
Finally, we will formulate a conjectural reconstruction algorithm for oriented matroids from VG algebras (4.5, Remark 5.5).
@@ 2.1. Hyperplane arrangements and oriented matroids
$\operatorname{Sep}(C,C^{\prime})=\{H\in\mathcal{A}\mid\text{$H$ separates $C$ and $C^{\prime}$}\}.$ (2.1)
@@
$E=\{\{C,C^{\prime}\}\mid\#\operatorname{Sep}(C,C^{\prime})=1\}.$ (2.2)
@@
$\operatorname{\mathcal{V}^{*}}(\bm{\alpha})=\{(\operatorname{sgn}\alpha_{1}(v),\dots,\operatorname{sgn}\alpha_{n}(v))\in\{+,-,0\}^{n}\mid v\in V\},$ (2.3)
@@
$\operatorname{\mathcal{V}}(\bm{\alpha})=\{(\operatorname{sgn}\lambda_{1},\dots,\operatorname{sgn}\lambda_{n})\in\{+,-,0\}^{n}\mid\lambda_{1}\alpha_{1}+\dots+\lambda_{n}\alpha_{n}=0,(\lambda_{i}\in\mathbb{R})\}.$ (2.4)
@@
$\bigcap_{i\in\operatorname{Supp}^{+}(\bm{\sigma})}H_{i}^{+}\cap\bigcap_{i\in\operatorname{Supp}^{-}(\bm{\sigma})}H_{i}^{-}=\emptyset.$ (2.6)
@@
$\operatorname{\mathcal{V}^{*}}(\alpha_{1},\dots,\alpha_{n})=\operatorname{\mathcal{V}^{*}}(\varepsilon_{1}\beta_{1},\dots,\varepsilon_{n}\beta_{n}).$ (2.7)
@@ 2.2. Intersection lattice and Orlik-Solomon algebra
$L(\mathcal{A})=\left\{\left.\bigcap_{H\in\mathcal{B}}H\right|\mathcal{B}\subset\mathcal{A}\right\}$ (2.8)
@@
$\partial e_{i_{1},\dots,i_{k}}=\sum_{p=1}^{k}(-1)^{p-1}e_{i_{1}}\wedge\cdots\wedge\widehat{e_{i_{p}}}\wedge\cdots\wedge e_{i_{k}}.$ (2.9)
@@
$\operatorname{\mathsf{OS}}_{R}^{\bullet}(\mathcal{A})=\frac{\wedge E}{\langle\partial e_{S}\mid S\subset[n]\mbox{ is dependent}\rangle}.$ (2.10)
@@ 2.3. Filtered and graded Varchenko-Gelfand algebras
$\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))=$ $\left\{\left.\sum_{C\in S}1_{C}\right|S\subset\operatorname{ch}(\mathcal{A})\right\}$ (2.12)
@@
$x_{i}^{+}(C)=\begin{cases}1,&(C\subset H_{i}^{+})\\ 0,&(C\not\subset H_{i}^{+}).\end{cases}$ (2.13)
@@
$\operatorname{\mathsf{VG}}^{k}(\mathcal{A})=\operatorname{\mathsf{F}}^{k}/\operatorname{\mathsf{F}}^{k-1},\ \ \operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})=\bigoplus_{k\geq 0}\operatorname{\mathsf{VG}}^{k}(\mathcal{A}),$ (2.16)
@@
    $\prod\limits_{i\in\operatorname{Supp}^{+}(\bm{\sigma})}e_{i}\cdot\prod\limits_{i\in\operatorname{Supp}^{-}(\bm{\sigma})}(e_{i}-1)-\prod\limits_{i\in\operatorname{Supp}^{+}(\bm{\sigma})}(e_{i}-1)\cdot\prod\limits_{i\in\operatorname{Supp}^{-}(\bm{\sigma})}e_{i}.$ (2.17)
@@
$\lambda_{i_{1}}\alpha_{i_{1}}+\dots+\lambda_{i_{k}}\alpha_{i_{k}}=0.$ (2.18)
@@
$\sum_{p=1}^{k}\operatorname{sgn}(\lambda_{i_{p}})\cdot e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{p}\}},$ (2.19)
@@
$0\longrightarrow\operatorname{\mathcal{VG}}(\mathcal{A}^{\prime})\longrightarrow\operatorname{\mathcal{VG}}(\mathcal{A})\stackrel{{\scriptstyle\rho_{H}}}{{\longrightarrow}}\operatorname{\mathcal{VG}}(\mathcal{A}^{H}).$ (2.21)
@@
$0\longrightarrow\operatorname{\mathsf{F}}^{k}\operatorname{\mathcal{VG}}(\mathcal{A}^{\prime})\longrightarrow\operatorname{\mathsf{F}}^{k}\operatorname{\mathcal{VG}}(\mathcal{A})\stackrel{{\scriptstyle\rho_{H}}}{{\longrightarrow}}\operatorname{\mathsf{F}}^{k-1}\operatorname{\mathcal{VG}}(\mathcal{A}^{H})\longrightarrow 0$ (2.22)
@@
$0\longrightarrow\operatorname{\mathsf{VG}}^{k}(\mathcal{A}^{\prime})\longrightarrow\operatorname{\mathsf{VG}}^{k}(\mathcal{A})\stackrel{{\scriptstyle\rho_{H}}}{{\longrightarrow}}\operatorname{\mathsf{VG}}^{k-1}(\mathcal{A}^{H})\longrightarrow 0$ (2.23)
@@ 3.1. Recovering from the set of Heaviside functions
$\mathcal{H}(\mathcal{A})=\{x_{i}^{\pm}\mid i=1,\dots,n\}.$ (3.1)
@@
$\{x\in\mathcal{H}(\mathcal{A})\mid x\cdot 1_{C}\neq x\cdot 1_{C^{\prime}}\}=2\cdot\#\operatorname{Sep}(C,C^{\prime}).$ (3.2)
@@
By definition of Heaviside functions, chambers $C$ and $C^{\prime}$ are separated by $H_{i}$ if and only if $x_{i}^{\pm}\cdot 1_{C}\neq x_{i}^{\pm}1_{C^{\prime}}$. The result follows immediately. ∎
By Proposition 3.3, $C$ and $C^{\prime}$ are adjacent if and only if $\{x\in\mathcal{H}(\mathcal{A})\mid x\cdot 1_{C}\neq x\cdot 1_{C^{\prime}}\}=2$. Therefore, once the set of Heaviside functions $\mathcal{H}(\mathcal{A})\subset\operatorname{\mathcal{VG}}(\mathcal{A})$ is known, the adjacency relation on the vertex set $\operatorname{PrimIdem}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ can be recovered algebraically. More precisely, the edge set can be constructed as
$\{\{1_{C},1_{C}^{\prime}\}\subset\operatorname{PrimIdem}(\mathcal{A})\mid\#\{x\in\mathcal{H}(\mathcal{A})\mid x\cdot 1_{C}\neq x\cdot 1_{C^{\prime}}\}=2\}$ (3.3)
@@
$\mathcal{H}(\mathcal{A})\subseteq\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\}.$ (3.4)
@@
In this situation, we can not algebraically distinguish $x_{1}^{\pm},x_{2}^{\pm},x_{3}^{\pm}$ from $y^{\pm}$. Indeed, the map $x_{1}^{+}\mapsto y^{-},x_{2}^{+}\mapsto x_{2}^{+},x_{3}^{+}\mapsto x_{3}^{+}$ defines an automorphism of the filtered algebra $\operatorname{\mathcal{VG}}(\mathcal{A})$. In this example, the automorphism group $\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ (see Section 3.3), of filtered algebra is isomorphic to the wreath product $\{\pm 1\}\wr\mathfrak{S}_{3}\simeq\{\pm 1\}^{3}\rtimes\mathfrak{S}_{3}$, a group of order $48$, which is strictly larger than the automorphism group of the tope graph (a group of order $12$).
@@ 3.2. Arrangements generic in codimension 22
$\operatorname{Supp}(y)=\{H\in\mathcal{A}\mid\rho_{H}(y)\neq 0\}.$ (3.6)
@@
$\#\mathcal{H}(\mathcal{A}_{1})$ $=\#\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}_{1}))\setminus\{0,1\}$ (3.8)
@@ 3.3. Automorphism groups
$\operatorname{Aut}_{\operatorname{\mathsf{alg}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))\simeq\operatorname{Aut}_{\operatorname{\mathsf{set}}}(\operatorname{ch}(\mathcal{A})),$ (3.9)
@@
    $\operatorname{Aut}_{\operatorname{\mathsf{graph}}}(\operatorname{\mathcal{T}}(\mathcal{A}))\subseteq\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))\subseteq\operatorname{Aut}_{\operatorname{\mathsf{set}}}(\operatorname{ch}(\mathcal{A})).$ (3.10)
@@
    $\operatorname{Aut}_{\operatorname{\mathsf{graph}}}(\operatorname{\mathcal{T}}(\mathcal{A}))=\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A})).$ (3.11)
@@
$\{C\in\operatorname{ch}(\mathcal{A})\mid C\subset H^{+}\}$ $=\{C\in\operatorname{ch}(\mathcal{A})\mid d(C,C_{0})<d(C,C_{1})\}$ (3.12)
@@
$\operatorname{Aut}_{\operatorname{\mathsf{graph}}}(\operatorname{\mathcal{T}}(\mathcal{A}))\subsetneq\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A})).$ (3.13)

### v2

@@ 1. Introduction
(Conjecture 4.3) Let $R$ be an integral domain with $\operatorname{char}R\neq 2$, and let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be real arrangements. If $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})_{R}\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})_{R}$ as graded $R$-algebras, then the oriented matroids of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are isomorphic.
@@
These results show that VG algebras contain sufficient information to recover oriented matroids under genericity conditions. We also point out that none of the following invariants determine the VG algebra: the intersection lattice $L(\mathcal{A})$, the Orlik-Solomon algebra $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A})$, or the Poincaré polynomial of $M(\mathcal{A})$ (Example 3.10, Example 5.2, Example 5.4). These relationships are summarized in Figure 1. (See also §2.4 for details.)
@@
Finally, we will formulate a conjectural reconstruction algorithm for oriented matroids from VG algebras (Conjecture 4.5, Remark 5.5).
@@ 2.1. Hyperplane arrangements and oriented matroids
(2.1) $\operatorname{Sep}(C,C^{\prime})=\{H\in\mathcal{A}\mid\text{$H$ separates $C$ and $C^{\prime}$}\}.$
@@
(2.2) $E=\{\{C,C^{\prime}\}\mid\#\operatorname{Sep}(C,C^{\prime})=1\}.$
@@
(2.3) $\operatorname{\mathcal{V}^{*}}(\bm{\alpha})=\{(\operatorname{sgn}\alpha_{1}(v),\dots,\operatorname{sgn}\alpha_{n}(v))\in\{+,-,0\}^{n}\mid v\in V\},$
@@
(2.4) $\operatorname{\mathcal{V}}(\bm{\alpha})=\{(\operatorname{sgn}\lambda_{1},\dots,\operatorname{sgn}\lambda_{n})\in\{+,-,0\}^{n}\mid\lambda_{1}\alpha_{1}+\dots+\lambda_{n}\alpha_{n}=0,(\lambda_{i}\in\mathbb{R})\}.$
@@
(2.6) $\bigcap_{i\in\operatorname{Supp}^{+}(\bm{\sigma})}H_{i}^{+}\cap\bigcap_{i\in\operatorname{Supp}^{-}(\bm{\sigma})}H_{i}^{-}=\emptyset.$
@@
(2.7) $\operatorname{\mathcal{V}^{*}}(\alpha_{1},\dots,\alpha_{n})=\operatorname{\mathcal{V}^{*}}(\varepsilon_{1}\beta_{1},\dots,\varepsilon_{n}\beta_{n}).$
@@ 2.2. Intersection lattice and Orlik-Solomon algebra
(2.8) $L(\mathcal{A})=\left\{\left.\bigcap_{H\in\mathcal{B}}H\right|\mathcal{B}\subset\mathcal{A}\right\}$
@@
(2.9) $\partial e_{i_{1},\dots,i_{k}}=\sum_{p=1}^{k}(-1)^{p-1}e_{i_{1}}\wedge\cdots\wedge\widehat{e_{i_{p}}}\wedge\cdots\wedge e_{i_{k}}.$
@@
(2.10) $\operatorname{\mathsf{OS}}_{R}^{\bullet}(\mathcal{A})=\frac{\wedge E}{\langle\partial e_{S}\mid S\subset[n]\mbox{ is dependent}\rangle}.$
@@ 2.3. Filtered and graded Varchenko-Gelfand algebras
(2.12) $\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))=$ $\left\{\left.\sum_{C\in S}1_{C}\right|S\subset\operatorname{ch}(\mathcal{A})\right\}$
@@
(2.13) $x_{i}^{+}(C)=\begin{cases}1,&(C\subset H_{i}^{+})\\ 0,&(C\not\subset H_{i}^{+}).\end{cases}$
@@
(2.16) $\operatorname{\mathsf{VG}}^{k}(\mathcal{A})=\operatorname{\mathsf{F}}^{k}/\operatorname{\mathsf{F}}^{k-1},\ \ \operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})=\bigoplus_{k\geq 0}\operatorname{\mathsf{VG}}^{k}(\mathcal{A}),$
@@
    (2.17) $\prod\limits_{i\in\operatorname{Supp}^{+}(\bm{\sigma})}e_{i}\cdot\prod\limits_{i\in\operatorname{Supp}^{-}(\bm{\sigma})}(e_{i}-1)-\prod\limits_{i\in\operatorname{Supp}^{+}(\bm{\sigma})}(e_{i}-1)\cdot\prod\limits_{i\in\operatorname{Supp}^{-}(\bm{\sigma})}e_{i}.$
@@
(2.18) $\lambda_{i_{1}}\alpha_{i_{1}}+\dots+\lambda_{i_{k}}\alpha_{i_{k}}=0.$
@@
(2.19) $\sum_{p=1}^{k}\operatorname{sgn}(\lambda_{i_{p}})\cdot e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{p}\}},$
@@
(2.21) $0\longrightarrow\operatorname{\mathcal{VG}}(\mathcal{A}^{\prime})\longrightarrow\operatorname{\mathcal{VG}}(\mathcal{A})\stackrel{{\scriptstyle\rho_{H}}}{{\longrightarrow}}\operatorname{\mathcal{VG}}(\mathcal{A}^{H}).$
@@
(2.22) $0\longrightarrow\operatorname{\mathsf{F}}^{k}\operatorname{\mathcal{VG}}(\mathcal{A}^{\prime})\longrightarrow\operatorname{\mathsf{F}}^{k}\operatorname{\mathcal{VG}}(\mathcal{A})\stackrel{{\scriptstyle\rho_{H}}}{{\longrightarrow}}\operatorname{\mathsf{F}}^{k-1}\operatorname{\mathcal{VG}}(\mathcal{A}^{H})\longrightarrow 0$
@@
(2.23) $0\longrightarrow\operatorname{\mathsf{VG}}^{k}(\mathcal{A}^{\prime})\longrightarrow\operatorname{\mathsf{VG}}^{k}(\mathcal{A})\stackrel{{\scriptstyle\rho_{H}}}{{\longrightarrow}}\operatorname{\mathsf{VG}}^{k-1}(\mathcal{A}^{H})\longrightarrow 0$
@@ 3.1. Recovering from the set of Heaviside functions
(3.1) $\mathcal{H}(\mathcal{A})=\{x_{i}^{\pm}\mid i=1,\dots,n\}.$
@@
(3.2) $\left\{x\in\mathcal{H}(\mathcal{A})\mid x\cdot 1_{C}\neq x\cdot 1_{C^{\prime}}\text{ and }0\in\{x\cdot 1_{C},x\cdot 1_{C^{\prime}}\}\right\}=2\cdot\#\operatorname{Sep}(C,C^{\prime}).$
@@
By definition of Heaviside functions, chambers $C$ and $C^{\prime}$ are separated by $H_{i}$ if and only if one of $x_{i}^{\pm}\cdot 1_{C}$ and $x_{i}^{\pm}\cdot 1_{C^{\prime}}$ is zero, and the other one is nonzero. The result follows immediately. ∎
By Proposition 3.3, $C$ and $C^{\prime}$ are adjacent if and only if $\{x\in\mathcal{H}(\mathcal{A})\mid x\cdot 1_{C}\neq x\cdot 1_{C^{\prime}},\text{ and }0\in\{x\cdot 1_{C},x\cdot 1_{C^{\prime}}\}\}=2$. Therefore, once the set of Heaviside functions $\mathcal{H}(\mathcal{A})\subset\operatorname{\mathcal{VG}}(\mathcal{A})$ is known, the adjacency relation on the vertex set $\operatorname{PrimIdem}(\operatorname{\mathcal{VG}}(\mathcal{A}))$, hence the tope graph $\operatorname{\mathcal{T}}(\mathcal{A})$, can be recovered algebraically.
@@
(3.3) $\mathcal{H}(\mathcal{A})\subseteq\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\}.$
@@
In this situation, we can not algebraically distinguish $x_{1}^{\pm},x_{2}^{\pm},x_{3}^{\pm}$ from $y^{\pm}$. Indeed, the map $x_{1}^{+}\mapsto y^{-},x_{2}^{+}\mapsto x_{2}^{+},x_{3}^{+}\mapsto x_{3}^{+}$ defines an automorphism of the filtered algebra $\operatorname{\mathcal{VG}}(\mathcal{A})$. In this example, the automorphism group $\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ (see §3.3), of filtered algebra is isomorphic to the wreath product $\{\pm 1\}\wr\mathfrak{S}_{3}\simeq\{\pm 1\}^{3}\rtimes\mathfrak{S}_{3}$, a group of order $48$, which is strictly larger than the automorphism group of the tope graph (a group of order $12$).
@@ 3.2. Arrangements generic in codimension 22
(3.5) $\operatorname{Supp}(y)=\{H\in\mathcal{A}\mid\rho_{H}(y)\neq 0\}.$
@@
(3.7) $\#\mathcal{H}(\mathcal{A}_{1})$ $=\#\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}_{1}))\setminus\{0,1\}$
@@ 3.3. Automorphism groups
(3.8) $\operatorname{Aut}_{\operatorname{\mathsf{alg}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))\simeq\operatorname{Aut}_{\operatorname{\mathsf{set}}}(\operatorname{ch}(\mathcal{A})),$
@@
    (3.9) $\operatorname{Aut}_{\operatorname{\mathsf{graph}}}(\operatorname{\mathcal{T}}(\mathcal{A}))\subseteq\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))\subseteq\operatorname{Aut}_{\operatorname{\mathsf{set}}}(\operatorname{ch}(\mathcal{A})).$
@@
    (3.10) $\operatorname{Aut}_{\operatorname{\mathsf{graph}}}(\operatorname{\mathcal{T}}(\mathcal{A}))=\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A})).$
@@
(3.11) $\{C\in\operatorname{ch}(\mathcal{A})\mid C\subset H^{+}\}$ $=\{C\in\operatorname{ch}(\mathcal{A})\mid d(C,C_{0})<d(C,C_{1})\}$
@@
(3.12) $\operatorname{Aut}_{\operatorname{\mathsf{graph}}}(\operatorname{\mathcal{T}}(\mathcal{A}))\subsetneq\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A})).$